OpenStudy (dls):

The radius of a circle is increasing at the rate of 0.7 cm/s.What is the rate of increase of its circumference?

OpenStudy (dls):

\[\LARGE C= 2 \pi R \] \[\LARGE \frac{dC}{dt}=2 \pi \frac{dR}{dt}=2 \pi \times 0.7\] is this correct?

OpenStudy (anonymous):

2pi*dr/dt

OpenStudy (anonymous):

looks like ur right

OpenStudy (anonymous):

to be exact

OpenStudy (anonymous):

1.4pi ceminiters

OpenStudy (anonymous):

or 2pi(0.7)

OpenStudy (dls):

but the solution says \[\LARGE \frac{dC}{dt}=\frac{dC}{dR} \times \frac{dR}{dt}\] why is it so? though answer is same

OpenStudy (anonymous):

how do i post pictures?

OpenStudy (dls):

attach file/draw

OpenStudy (anonymous):

That's just the chain rule. Are you asking for a proof of the rule, or why they wrote the answer that way?

OpenStudy (dls):

latter!

OpenStudy (dls):

i mean simply what i wrote in first line is wrong?

OpenStudy (anonymous):

No, it's not wrong. Their statement is more general, yours is specific to the question, but they're both right.

OpenStudy (anonymous):

The circumference of a circle (C) with radius (r) is given by C = 2πr. Therefore, the rate of change of circumference (C) with respect to time (t) is given by,

OpenStudy (anonymous):

ur welcome

OpenStudy (dls):

i got it now :D thanks